# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

### v1

@@ 5.1 Fluid-dynamic model
The gas flow dynamic during injection and withdrawal is simulated by different codes according to the selected fluid: OPM Flow (The Open Porous Media Initiative, 2023; Rasmussen et al., 2021) for CH4 (UGS), and Eclipse (Schlumberger, 2014) for H2 (UHS), CO2 (CCS), and N2. For all simulations, the same well configuration is used. During the PP stage, two wells located next to the reservoir boundary are open in the first layer. Then, these wells are converted to injection purposes. Only for UHS, nine storage wells serve for injection and withdrawal. These wells are evenly distributed throughout the reservoir. In Figure 5, the location of the injection and production wells is shown with respect to the fault system. To avoid interpolation between different computational grids, both OPM and Eclipse finite volume cells coincide with the hexahedral elements of the geomechanical model. Each reservoir block is subdivided into 4000 regular cells with a $20\times 20\times 10$ partitioning. The characteristic horizontal and vertical permeability is $k_{h}=600$ mD and $k_{v}=300$ mD, respectively.
Figure 5: Areal location of the injection/production wells (blue and red for injection/production for CH4, CO2, N2 and PP for H2; green for UHS cycles) and example of pressure distribution from a dynamic simulation for CH4 (on top) and H2 on bottom within a reservoir block. The grid is split to provide evidence of the pressure distribution along the vertical direction.
Figure 6: Maximum pore pressure variation in time for the investigated fluids. The leftmost panel is the common pressure decline during the primary production stage.
The injection-production history is depicted in Figure 6. The PP stage is conventionally set to 10 years with a maximum pore pressure drop of about 18 MPa. The targeted final Recovery Factor, i.e., the ratio between the cumulative produced volume of gas and the gas originally in place, is 90%, in line with theoretical reference behavior for a depletion-driven gas reservoir. Then, for CH4 the reservoir pressure recovers to the initial value in two years during the CGI phase and a set of UGS cycles, consisting of 6-month production and 6-month storage each with a 10-MPa pressure excursion, follows. The simulation results show that the pressure change throughout the production (or injection) phase in a single UGS cycle remains nearly uniform in space. Hence, employing a uniform pressure change within each block at every loading step could be also fully warranted. As to the other fluids, the pressure history is targeted to closely approximate the maximum drop in reservoir static pressure observed for CH4, with some adjustments needed to achieve more realistic storage strategies. In particular, the 20-MPa pressure recovery increment during CGI is carried out in a longer period and no production/injection cycles are simulated for CO2 and N2 storage. The CO2 storage scenario considers a maximum injection rate of 4,106 Sm3/day per well for 13 years before recovering the initial pressure value. Injection of N2 lasts 6 years with an injection rate of 4,106 Sm3/day per well. Concerning H2, 10 years of injection/production activity are simulated after a 13-year refilling phase at a maximum rate of 4,106 Sm3/day per well. Each UHS cycle consists of a 6-month withdrawal and a 6-month injection period with a maximum pressure excursion of 9.2 MPa. In all simulations, a 1-year loading step (l.s.) is used during PP and CGI/refilling, while during UGS and UHS the loading step is equal to 15 days.
It is important to recall that in the Netherlands the fluid pressure is not allowed to exceed the natural initial value $P_{i}$. Pore pressure propagation is also accounted for by simulating its variation within a permeable finite-thickness fault zone. The fluid pressure in the fault, $P_{f}$, generates forces that opposes to the action of the pressure variation on $\sigma_{n}$.
@@ 5.2 Mechanisms of fault re-activation: reference case (Scenario #1)
The basic fault reactivation mechanisms are studied in scenario #1, which serves as the reference scenario for UGS activity with CH4. A few preliminary outcomes can already be found in Teatini et al. (2019, 2020) and Franceschini et al. (2024). For the sake of completeness, they are briefly summarized here as well.
@@
Figure 7 shows the behavior of the maximum $\chi$ value, $\chi_{max}$, experienced on faults F1 through F5 for 13 years, i.e., PP, CGI and the first UGS cycle. Faults F1 and F2 can reach a critical condition at the end of the PP stage. During CGI and UGS, the fault activation risk reduces, with $\chi_{max}$ decreasing to about 0.80. During CGI, $\chi_{max}$ initially decreases, i.e., the fault is more stable, but then increases again, with the shear stress acting in the opposite direction with respect to that experienced during PP. The time behavior of the maximum sliding $\delta_{max}$ computed for each fault is also shown in Figure 7. These two quantities, i.e., $\chi_{max}$ and $\delta_{max}$, are closely related each other, since a single element can slide only when $\chi=1$. However, we prefer to propose an averaged version of $\chi$, so as to obtain information on the criticality state of the entire fracture at a given depth. In particular, for the sake of clarity and ease of readability, $\chi$ is represented for each fault as a function of depth only, i.e., for each $z$-value we compute the $\chi$ average for the stripe of fault elements located at the same depth. For this reason, it is possible to detect sliding conditions even when the $\chi$ value reported on the top frame of Figure 7 is smaller than 1. Fault F1, F2, F4, and F5 can start sliding between year 4 and 6. The maximum sliding can be of about 1.4 cm. Fault F3 remains unaffected throughout the loading and injection steps due to the model symmetry. The most critical condition develops along the top and bottom of the reservoir. Figure 8 shows the behavior of $\chi$ with respect to depth on all the fracture surfaces at loading step 10.
@@
Figure 7: Scenario #1: $\chi_{max}$ (top), $\delta_{max}$ (middle), and $\tau$ in time for all faults. Positive values for the shear stress means that it is upward oriented.
Figure 8: Scenario #1: $\chi$ in depth at the PP end taken at the red point (on the left sketch).
@@
The mechanisms leading to the possible fault reactivation can be better understood in terms of shear stress $\tau$ (bottom frame of Figure 7), particularly at at loading steps 0, 10 (end of PP), 11, 12 (end of CGI), 12.5 (end of UGS extraction), and 13 (end of UGS injection). The largest values are computed at the end of primary production, with opposing shear stress orientations at top and bottom of the reservoir. During PP, the direction of the shear stress is always oriented toward the reservoir mid-plane. At loading step 4 and 5 sliding starts (middle of Figure 7) at reservoir top and bottom. The elements surrounding the activated fault portions increase their shear stress to accommodate the excess to the Coulomb limit not supported by the activated (sliding) elements. At loading step 11, half of the pore pressure change has been recovered. As the reservoir expands due to pressure recovery, the shear stress changes direction and annihilates approximately the shear stress acting on the previously sliding elements. The reservoir continues to recover pressure and re-expand until loading step 12. During the second half of cushion gas injection the shear stress increases again, with a direction opposite to that experienced during primary production. The expansion during CGI can generate again a critical condition on the fault, with a possible localized reactivation at the reservoir top due the stress redistribution following the sliding. During UGS, at the end of the production phase (loading step 12.5), the shear stress almost equals the condition at loading step 11. Correspondingly, the shear stress at loading step 13 equals the stress state at loading step 12. Indeed, during UGS, sliding does not occur and the porous medium behaves according to a linear elastic constitutive law. This complex behavior is also confirmed by the analysis of the stress path experienced by some representative elements, as reported by Franceschini et al. (2024).
@@ 5.3 Sensitivity analysis: Stage 1
Starting from the reference configuration of scenario #1, we explore how variations in individual geometric or mechanical parameters can affect the outcomes. We conducted an extensive investigation, of which we report here the most significant results referring the scenarios summarized in Table 2 for the UGS analysis with CH4. Other simulations were carried out, but are not discussed here for the sake of brevity. The outcomes are cross compared to identify the settings that make the subsurface system more prone to fault reactivation. For the analysis, we use the parameter $t_{80}$, i.e., the areal extent where $\chi\geq 0.80$. Such areal extent is scaled by 2,000 m, which is the characteristic size of each reservoir block, so that $t_{80}$ has the size of a length. For the sake of simplicity, we focus on faults F2 and F3, unless differently specified.
@@ 5.3.1 Reservoir and fault geometry
The impact of fault geometry is investigated by varying the dip angle of the central fault F3 (#2a - #2b) and increasing the offset between reservoir compartments of half (#2c) and entire (#2d) reservoir thickness. Figure 9 shows $\chi_{max}$ in time for all faults in all scenarios. Interestingly, while the geometry variation has no effect on fault F1, a large offset between the reservoir compartments (scenario #2d) enhances the potential for instability in faults F4 and F5. The instability of fault F2 is slightly increased by a variation of the fault geometry, however this becomes significant only when a large offset between compartments is introduced (#2d).
Figure 9: Scenarios #2: behavior of $\chi_{max}$ from all the loading steps for each fault. Note that F4 and F5 exhibit an identical behavior, due to symmetric conditions.
Figure 10: Scenario #2d: maximum sliding $\delta_{max}$.
In the latter configuration, unlike the reference scenario, fault F3 is significantly affected by the asymmetry of the system. Indeed, it becomes active and experiences the highest stress levels at the top and bottom of each compartment. Unlike in the reference scenario, now $\chi_{max}$ on fault F3 can reach the limiting value. Furthermore, $\chi_{max}$ achieves the alert threshold of 0.8 also during CGI and UGS, suggesting that UGS operations in this condition might be critical. Consistently, fault F3 exhibits the highest sliding value $\delta_{max}$ of 7.4 cm at the end of the PP stage (Figure 10). Some instability can also occur at the end of the UGS cycle. The other faults can activate around the loading step 4 again at the reservoir top first, and at the reservoir bottom later. Figure 11 shows the active elements within the embedded fault discretization at loading step 6, 10 (end of PP) and 12.5 (middle of a UGS cycle).
@@ 5.3.2 Geo-mechanical parameters of faults and reservoir
$\varphi=\begin{cases}\varphi_{s}+\dfrac{\varphi_{d}-\varphi_{s}}{d_{c}}\delta&\text{for }\delta<d_{c}\\ \varphi_{d}&\text{for }\delta>d_{c}\end{cases}$ (2)
The two parameters defining the new constitutive law are $\varphi_{d}$ and $d_{c}$, i.e., the dynamic friction angle and the slip weakening distance, respectively. Slip weakening occurs after sliding begins by reducing the value of the friction angle. The most critical results in terms of fault stability are obtained for scenario #3c, characterized by a reduction of the friction angle from $\varphi_{s}=30^{\circ}$ to $\varphi_{d}=10^{\circ}$ in a slip distance of $d_{c}=2$ mm. The behavior of $\chi_{max}$ for scenarios #3 during the CGI and UGS stages is shown in Figure 12, showing that slip-weakening may cause a reactivation also at loading steps 11 and 13, but not at loading step 12.5, i.e., in the middle of a UGS cycle.
Figure 12: Scenarios #3: $\chi_{max}$ in time over the CGI and UGS phases.
Figure 13: Scenarios #4: $\chi_{max}$ in time over the CGI and UGS phases.
The relationship between reservoir stiffness and fault behavior has been further investigated in Scenarios #4. A significant contrast between the reservoir and the surrounding caprock, sideburden, and underburden stiffness may play an important role to induce a critical stress state on faults. In scenario #6a the reservoir is softer than the surroundings, and this generally increases the likelihood of a fault reactivation (see Figure 13). Conversely, in scenario #4b the fault reactivation likelihood is notably reduced. During the UGS stage, $\chi_{max}$, remains below 0.7, regardless of the reservoir stiffness value (see Figure 13).
@@ 5.3.3 Uneven pore pressure excursion in the reservoir compartments
The effect of a different pore pressure change in the two reservoir compartments is investigated in Scenarios #5 (Table 2). During UGS, fault F3 and, secondarily, faults F4 and F5, may exhibit larger $\chi$ values at loading step 12.5 (Figure 14). When $\Delta P_{2}=0$ MPa, faults F4 and F5 keeps the same critical degree experienced at the end of CGI over the UGS phase. Conversely, with $\Delta P_{2}=20$ MPa the most critical condition is achieved in the middle of the UGS cycle, where $\chi_{max}$ is larger than the threshold value 0.80. Figure 15 shows the distribution of the potentially active elements ($\chi\geq 0.80$) in Scenarios #5, again along the reservoir boundaries for block 2.
Figure 14: Scenarios #5: $\chi_{max}$ in time over the CGI and UGS phases.
@@
For scenario #5b, Figure 16 shows the time behavior of $\chi_{max}$ and the shear stress modulus $\tau$ for a node located at the reservoir bottom on fault F4. The shear stress increases during the PP stage, reaching the limit shear stress $\tau_{L}$ at loading step 5. Then, $\tau$ equates $\tau_{L}$ at the end of CGI and at the end of UGS production and injection cycle. On fault F2 $\chi_{max}$ approaches 1 at the end of the UGS injection phase too, but remains well below the criticality value at the end of CGI phase.
Figure 16: Scenario #5b: $\chi_{max}$ in time on fault F1 (left), $\|\boldsymbol{\tau}\|_{2}$ and $\tau_{L}$ in time on fault F4 (right).

### v4

@@ 3.1 Fluid-dynamic model
The results of the fluid-dynamic model indicate that, during CH4 storage, the pressure change within each UGS cycle remains nearly uniform across the reservoir block during both withdrawal and injection phases. This behavior is consistent with the spatial pressure distributions shown in Figure 5 c,d, where the normalized pressure fields for CH4 and H2 exhibit limited lateral variations within the same block. The vertical cuts displayed in Figure 5 c,d further confirm that the pressure variations are essentially homogeneous across the block thickness.
Figure 6 illustrates the imposed pore-pressure evolution for all investigated fluids. For CH4, the pressure returns to the initial value following the PP stage and the subsequent CGI phase, after which the prescribed UGS cycles generate the imposed 10-MPa oscillations (Figure 6 a,b) . For CO2 and N2, the pressure increases according to the imposed injection rates until the initial pressure is recovered (Figure 6 c,e). For H2, the imposed refilling and cycling schedule leads to the 10-MPa pressure variations associated with the UHS cycles (Figure 6 d).
Figure 6: Schematic evolution of pore-pressure variation for the investigated fluids.(a) Primary production ($P$): 10-year pressure decline of about 20 MPa. (b) CH4: 2-year cushion-gas injection (CGI) followed by UGS cycles consisting of 6-month withdrawal and 6-month injection (10 MPa swing). (c) CO2: 10-year pressure recovery toward $P_{i}$. (d) H2: 8-year refilling phase up to $P_{i}$, followed by 6-month withdrawal and 6-month injection (10 MPa swing) UHS cycles (e) N2: 10-year injection stage reaching $P_{i}$.
@@ 3.2 Mechanisms of fault re-activation: reference case (Scenario #1)
The basic fault reactivation mechanisms are studied in Scenario #1, which serves as the reference scenario for UGS activity with CH4. Scenario #1 does not represent the full structural complexity of real Rotliegend reservoirs; rather, it is a simplified setting adopted to clarify the fundamental mechanical processes governing fault reactivation during PP, CGI and UGS. The choice of this configuration was made in consultation with Staatstoezicht op de Mijnen (SodM) , and aims to reproduce the essential geological, stress and operational conditions of a typical Rotliegend UGS reservoir while maintaining a geometry that allows the isolation of the primary mechanical mechanisms before exploring more complex configurations. For this reason, Scenario #1 serves as the "reference" against which all subsequent parametric variations are interpreted. A few preliminary outcomes can already be found in previous studies (Teatini et al., 2019, 2020; Franceschini et al., 2024). For the sake of completeness, they are briefly summarized here as well.
@@
Figure 7 a shows the behavior of the maximum $\chi$ value, $\chi_{max}$, experienced on faults F1 through F5 for 13 years, i.e., PP, CGI and the first UGS cycle. Faults F1 and F2 can reach a critical condition at the end of the PP stage. During CGI and UGS, the fault activation risk reduces, with $\chi_{max}$ decreasing to about 0.80. During CGI, $\chi_{max}$ initially decreases, i.e., the fault returns more stable, but then increases again, with the shear stress acting in the opposite direction with respect to that experienced during PP (Figure 7 c - orange panel). The time behavior of the maximum sliding $d_{max}$ computed for each fault is also shown in Figure 7 b. These two quantities, i.e., $\chi_{max}$ and $d_{max}$, are closely related to each other, since a single element can slide only when $\chi=1$. Notice that, to obtain information on the criticality state of the entire fracture at a given depth, the provided $\chi_{max}$ represents the value averaged on the stripe of fault elements located at the same depth. For this reason, it is possible to detect sliding conditions even when the $\chi$ value reported on Figure 7 a is smaller than 1. Fault F1, F2, F4, and F5 can start sliding between year 4 and 6. The maximum sliding can be of about 1.4 cm. Fault F3 remains unaffected throughout the loading and injection steps due to the model symmetry. This behavior, however, directly reflects the simplified structural setting adopted in Scenario #1. In this configuration, fault F3 is perfectly vertical and exhibits no offset, so no shear stress builds up during PP or CGI/UGS. As a consequence, the bounding faults (F1–F2 and F4–F5) display the largest shear variations, since they juxtapose a depleting reservoir block against sideburden regions.
The behaviour of $\chi$ with respect to depth on all the fracture surfaces at loading step 10 is shown in Figure 24 in the Supplementary Materials. These profiles indicate that the most critical conditions develop along the top and bottom of the reservoir.
@@
Figure 7: Scenario #1 ("Reference"): $\chi_{max}$ (a) and $d_{max}$ (b) over time for all faults. (c) behavior of $\tau$ for three nodes located at the top, bottom, and center of F1. Positive values for the shear stress means that it is upward oriented. These plots summarize the reactivation mechanisms during PP, CGI and UGS in the simplified reference configuration.
@@
The mechanisms leading to the possible fault reactivation can be better understood in terms of shear stress $\tau$ (7 c), particularly at loading steps 0, 10 (end of PP), 11, 12 (end of CGI), 12.5 (end of UGS withdrawal), and 13 (end of UGS injection). The largest values are computed at the end of PP, with opposing shear stress orientations at top and bottom of the reservoir. During PP, the direction of the shear stress is always oriented towards the reservoir mid-plane. At loading step 4 and 5 sliding starts (Figure 7 b) at reservoir top and bottom. The elements surrounding the activated fault portions increase their shear stress to accommodate the excess to the Coulomb frictional limit not supported by the activated (sliding) elements. At loading step 11, half of the pore pressure change has been recovered. As the reservoir expands due to pressure recovery, the shear stress changes direction and annihilates approximately the shear stress acting on the previously sliding elements. The reservoir continues to recover pressure and re-expand until loading step 12. During the second half of CGI, the shear stress increases again, with a direction opposite to that experienced during PP. The expansion during CGI can generate again a critical condition on the fault, with a possible localized reactivation at the reservoir top due to the stress redistribution following the sliding. During UGS, at the end of the production phase (loading step 12.5), the shear stress almost equals the condition at loading step 11. Correspondingly, the shear stress at loading step 13 equals the stress state at loading step 12. Indeed, during UGS, no new sliding occurs. The faults keep the slip accumulated during PP and CGI, and the porous medium behaves according to a linear elastic constitutive law.
These results, together with the stress-path interpretation developed in Franceschini et al. (2024), provide the basis for understanding the occurrence of “unexpected” reactivation during CGI or UGS, even though the reservoir pressure remains below $p_{i}$. The slip accumulated during primary production produces a permanent redistribution of stresses along the fault system. As a result, when pressure increases during CGI and UGS, the response does not follow the elastic unloading path of PP. Instead, the system evolves along a different stress trajectory, characterized by a reversal of the shear stress $\tau$ at the top and bottom of the reservoir (see Figure 7 c). Because of this stress-path reversal, the fault may approach the failure condition again even at pressures within the previously experienced range. The renewed criticality does not require exceeding the historical maximum loading, but it arises because the system returns toward failure from a different direction in the $(\sigma_{n},\tau)$ space, driven by the stress redistribution inherited from PP slip (Franceschini et al., 2024) To illustrate this mechanism in simple terms, we consider the evolution of $\chi_{max}$ on the bounding faults F1 and F2 as an example. During UGS, the imposed pressure variation is of the order of 10 MPa. If such a pressure cycle were applied to a pristine reservoir–fault system, the resulting criticality would remain moderate (e.g., $\chi_{max}\approx 0.6$), indicating stable conditions. In the modeled case, however, the same pressure variation acts on a fault system that has already experienced PP and partial pressure recovery. Slip occurring during PP on faults F1 and F2 produces a permanent redistribution of stresses, so that the subsequent CGI and UGS phases do not correspond to a simple elastic unloading–reloading of the initial state. As a result, the same $\Delta P$ applied during UGS leads to significantly higher criticality levels on these faults (e.g., $\chi_{max}\approx 0.8$), increasing the likelihood of approaching failure even though the stress magnitudes remain lower than those reached during PP.
@@ 3.3 Sensitivity analysis: Stage 1
Starting from the reference configuration of scenario #1, we explore how variations in individual geometric or mechanical parameters can affect the model outcomes, including fault reactivation potential and the associated stress and slip patterns. We conducted an extensive investigation, of which we report here the most significant results, referring to the scenarios summarized in Table 4 for the UGS analysis with CH4. Other simulations were carried out, but are not discussed here for the sake of brevity. The outcomes are cross-compared to identify the settings that make the subsurface system more prone to fault reactivation. For the analysis, we use the parameter $t_{80}$, i.e., the areal extent where $\chi\geq 0.80$. Such areal extent is scaled by 2,000 m, which is the characteristic size of each reservoir block, so that $t_{80}$ has the size of a length. For the sake of simplicity, we focus on faults F2 and F3, unless differently specified.
@@ 3.3.1 Reservoir and fault geometry
The impact of fault geometry is investigated by varying the dip angle of the central fault F3 (#2a - #2b) and increasing the offset between reservoir compartments of half (#2c) and entire (#2d) reservoir thickness. Figure 8 shows $\chi_{max}$ over time for the reference and the #2a - #2d scenarios. Interestingly, while the geometry variation has no effect on fault F1, a large offset between the reservoir compartments (scenario #2d) enhances the potential for instability in faults F4 and F5. The instability of fault F2 is slightly increased by a variation of the fault geometry, however this becomes significant only when a large offset between compartments is introduced (#2d).
Figure 8: Scenarios #2: $\chi_{max}$ over time for each fault. Note that F4 and F5 exhibit an identical behavior, due to symmetric conditions. Scenario 2 includes four geometrical variants: (#2a) F3 dip $65^{\circ}$, (#2b) F3 dip $65^{\circ}$, (#2c) Block 2 lowered by 100 m, (#2d) Block 2 lowered by 200 m.
Figure 9: Scenario #2d: maximum sliding $d_{max}$. Scenario #2 includes (#2a) F3 dip $65^{\circ}$, (#2b) F3 dip $65^{\circ}$, (#2c) Block 2 lowered by 100 m, (#2d) Block 2 lowered by 200 m.
In the latter configuration, unlike the reference scenario, fault F3 is significantly affected by the asymmetry of the system. Indeed, it becomes active and experiences the highest stress levels at the top and bottom of each compartment. Differently from the reference scenario, $\chi_{max}$ on fault F3 can now reach the limiting value. Furthermore, during CGI and UGS, $\chi_{max}$ reaches approximately 0.8. Although this does not correspond to reactivation, it indicates that the fault is close to a critically stressed state.
Furthermore, $\chi_{max}$ achieves the alert threshold of 0.8 also during CGI and UGS, suggesting that UGS operations in this condition might be critical. Consistently, fault F3 exhibits the highest sliding value $d_{max}=7.4$ cm at the end of the PP phase (Figure 9). Some instability can also occur at the end of the UGS cycle. The other faults can activate around the loading step 4, again at the reservoir top first, and at the reservoir bottom later. The active elements within the embedded fault discretization at loading step 6, 10 (end of PP), and 12.5 (middle of a UGS cycle) are shown in Figure 25 in the Supplementary Materials.
@@ 3.3.2 Geo-mechanical parameters of faults and reservoir
$\varphi=\begin{cases}\varphi_{s}+\dfrac{\varphi_{d}-\varphi_{s}}{d_{c}}d&\text{for }d<d_{c}\\ \varphi_{d}&\text{for }d\geq d_{c}\end{cases}$ (2)
The two parameters defining the new constitutive law are $\varphi_{d}$ and $d_{c}$, i.e., the dynamic friction angle and the slip weakening distance, respectively. Slip weakening occurs after sliding begins by reducing the value of the friction angle. The most critical results in terms of fault stability are obtained for scenario #3c, characterized by a reduction of the friction angle from $\varphi_{s}=30^{\circ}$ to $\varphi_{d}=10^{\circ}$ in a slip distance of $d_{c}=2$ mm. The behavior of $\chi_{max}$ for scenarios #3 during the CGI and UGS stages is reported in Figure 10, showing that slip-weakening may cause a reactivation also at loading steps 11 and 13, but not at loading step 12.5, i.e., in the middle of a UGS cycle.
Figure 10: Scenarios #3: $\chi_{max}$ over time for the CGI and UGS phases.
Scenario #3 includes four fault–strength variants: (#3a) cohesion $c=0.5$ MPa, (#3b) friction angle $\varphi_{s}=20^{\circ}$, (#3c) slip–weakening with $\varphi_{d}=10^{\circ}$ and $d_{c}=2$ mm, (#3d) slip–weakening with $\varphi_{d}=20^{\circ}$ and $d_{c}=20$ $m$.
Figure 11: Scenarios #4: $\chi_{max}$ over time for the CGI and UGS phases. Scenario #4 includes (#4a) Young’s modulus $E=8$ GPa, (#4b) Young’s modulus $E=20$ GPa.
The relationship between reservoir stiffness and fault behavior has been further investigated in Scenarios #4. A significant contrast between the reservoir and the surrounding caprock, sideburden, and underburden stiffness may play an important role to induce a critical stress state on faults. In scenario #4a, the reservoir is softer than the surroundings, and this generally increases the likelihood of a fault reactivation (see Figure 11). Conversely, in scenario #4b, the fault reactivation likelihood is notably reduced. During the UGS stage, $\chi_{max}$ remains below 0.7, regardless of the reservoir stiffness value (see Figure 11).
@@ 3.3.3 Uneven pore pressure excursion in the reservoir compartments
The effect of a different pore pressure change in the two reservoir compartments is investigated in Scenarios #5 (Table 4). During UGS, fault F3 and, secondarily, faults F4 and F5, may exhibit larger $\chi$ values at loading step 12.5 (Figure 12). When $\Delta P_{2}=0$ MPa, faults F4 and F5 keeps the same critical degree experienced at the end of CGI over the UGS phase. Conversely, with $\Delta P_{2}=-20$ MPa the most critical condition is achieved in the middle of the UGS cycle, where $\chi_{max}$ is larger than the threshold value 0.80. Figure 13 shows the distribution of the potentially active elements ($\chi\geq 0.80$) in Scenarios #5, again within the embedded fault discretization at loading step 12.5.
Figure 12: Scenarios #5: $\chi_{max}$ over time for the CGI and UGS phases. Scenario #5 includes two uneven–pressure variants: (#5a) $\Delta P_{1}=-10$ MPa and $\Delta P_{2}=0$ MPa, (#5b) $\Delta P_{1}=-10$ MPa and $\Delta P_{2}=-20$ MPa.
@@
For scenario #5b, Figure 14 (a) shows the time behavior of $\chi_{max}$ and (b) the shear stress modulus $\tau$ at the reservoir bottom on fault F4 . The shear stress increases during the PP stage, reaching the limit shear stress $\tau_{L}$ at loading step 5. Then, $\tau$ equates $\tau_{L}$ at the end of CGI and at the end of UGS production and injection cycle. On fault F2, $\chi_{max}$ approaches 1 at the end of the UGS injection phase too, but remains well below the criticality value at the end of CGI phase.
Figure 14: Scenario #5b: (a) $\chi_{max}$ over time on fault F1 , (b) temporal evolution of the shear stresses $\tau$ and $\tau_{L}$ on fault F4, showing how the shear loading evolves during PP, CGI and UGS.